OpenStudy (anonymous):

Find the equilibrium point for the demand and supply function D(p)=79-8p and S(p)=37+6p

OpenStudy (amistre64):

so, how do we determine when supply equals demand?

OpenStudy (amistre64):

D(p) = S(p) is the equilibrium point D(p) - S(p) = 0 ; and solve for p

OpenStudy (anonymous):

im confused. how would i find p?

OpenStudy (amistre64):

by subtracting the value of S(p) from D(p) .. i thought i wrote that?

OpenStudy (amistre64):

D(p)= 79-8p -S(p)=-(37 +6p) ---------------- 0 = 42 - 2p then solve for p

OpenStudy (anonymous):

the website crashed for a few minutes... that might have not been placed in time

OpenStudy (anonymous):

wouldnt it be 0=42 -14p?

OpenStudy (anonymous):

that would be 3 but would i put the answer as (3,0) ?

OpenStudy (amistre64):

you are correct; 42 -14p

OpenStudy (amistre64):

42/14 = 21/7 = 3 the market equilibrium price would be 3 (m,k,b, or other unit) price. the quantity supplied is then whatever we get when we plug "3" into the supply equation

OpenStudy (amistre64):

S(p) = 37+6p S(3) = 37+6(3) = 37+18 = 56

OpenStudy (amistre64):

(3,56) if that is the format they want, but of that i cant be sure